# Revisiting Data Challenges of Computational Pathology: A Pack-based Multiple Instance Learning Framework

### v1

@@ 3.3 Training Recipe
Task-specific Hyperslide Labels and Loss. Heterogeneous pathological data and diverse downstream clinical tasks necessitate specialized label generation strategies that preserve task-relevant clinical characteristics and enable efficient learning from multiple WSIs, encouraging the design of task-specific hyperslide labeling mechanisms. We divide downstream tasks into two classes: Morphological Categorical and Event-Driven. Based on pathological scenarios, we design three strategies to generate a hyperslide label $y^{\mathrm{hyper}}$ for higher-level supervision, illustrated in Fig.4. Appropriate loss functions are selected for different tasks, as detailed in the Appendix D.1.
1) Grading. Let each pack contain $S$ WSIs with pathologist-assigned scores or tumour-area ratios $g_{s}$ and corresponding patch counts $n_{s}$. We compute a pack-level statistic by $\tilde{g}=\frac{\sum_{s=1}^{S}n_{s}\,g_{s}}{\sum_{s=1}^{S}n_{s}}$. The weighted score (or ratio) is then mapped to the discrete grading rubric recommended by the pathology guideline, yielding a single categorical hyperslide label $y^{\mathrm{hyper}}\in\{1,\dots,G\}$. Weighting by $n_{s}$ ensures that the pack-level grading reflects the relative tissue coverage of each WSI, preventing smaller tissue regions from being overrepresented in the final assessment.
@@
where $\mathcal{E}$ is the ordered event set. This strategy ensures that each bag is uniquely associated with the most clinically significant event, aligning label semantics closely with clinically practice.
Practical Guidelines for Batched Training. Selecting an appropriate batchsize ($bs$) across various CPath tasks remains an underexplored yet critical problem for effective model training. To establish an batched training protocol, we conduct a series of preliminary experiments. The results from these experiments inform the practical guidelines discussed below, which are subsequently adopted in our main experimental setup. We find that the specifics of these guidelines vary significantly across benchmarks and are strongly correlated with dataset scale.
On conventional CPath datasets (e.g., TCGA), which typically contain a limited number of gigapixel WSIs, resource constraints often force a trade-off between the $bs$ and the number of patches sampled per WSI (NP), as shown in the right figure(b). The optimal choice of $bs$ usually requires empirical tuning. Additionally, regular rules for scaling the learning rate do not apply and also necessitate empirical tuning, as depicted in the right figure(c).
On large-scale CPath datasets (e.g., PANDA Bulten et al. (2022)), the $\sqrt{bs}$ learning rate scaling rule often proves effective. However, performance does not monotonically increase with $bs$; excessively large values can degrade performance, as illustrated in the right figure(a). Moreover, we found that 1D Batch Normalization Ioffe & Szegedy (2015) can effectively facilitate convergence on benchmarks with sufficient data scale. In contrast, normalization does not provide significant improvements on conventional CPath datasets, which we attribute to their limited data volume.
@@ 4.1 Datasets and Evaluation Metrics
For cancer diagnosis, we evaluate performance on cancer grading and sub-typing tasks using the PANDA Bulten et al. (2022) and TCGA-BRCA datasets. For cancer prognosis, we evaluate survival analysis performance using TCGA-LUAD and TCGA-BRCA. We report macro accuracy (Acc.) for cancer grading and the area under the ROC curve (AUC) for sub-typing. For survival analysis, we report the concordance index (C-index) Harrell Jr et al. (1996). To ensure statistical robustness, we perform 1000 bootstrap iterations and repeat experiments across multiple random seeds. We report the mean and 95% confidence interval for all metrics. Appendix A gives further details.
Table 1: Comparison of ISUP Grading (Acc.) on PANDA and Sub-typing (AUC) on BRCA.
@@
ABMIL Ilse et al. (2018) 65.48±0.70 73.21±0.61 72.91±0.62 89.58±5.15 93.58±3.92 94.13±3.88
DSMIL Li et al. (2021) 71.95±0.72 71.41±0.68 70.84±0.73 90.68±4.91 93.89±3.79 93.70±4.14
CLAM Lu et al. (2021) 66.91±0.73 74.76±0.61 74.99±0.60 89.61±5.05 93.90±3.86 93.73±3.75
TransMIL Shao et al. (2021) 68.53±0.79 72.59±0.75 71.91±0.74 91.91±4.34 94.09±3.79 93.57±3.84
DTFD Zhang et al. (2022) 63.57±0.71 72.19±0.66 71.93±0.65 91.07±4.59 93.86±3.93 93.88±3.81
WiKG Li et al. (2024c) 72.37±0.72 76.68±0.68 74.91±0.68 91.93±4.65 93.81±4.25 94.06±3.92
RRTMIL Tang et al. (2024) 69.99±0.69 72.18±0.67 72.34±0.63 91.67±4.77 94.10±3.65 94.04±3.86
2DMamba Zhang et al. (2024a) 71.59±0.73 74.97±0.68 75.36±0.66 90.96±3.85 93.59±3.14 93.33±3.25
ABMIL+RS 74.72 +9.2 77.91 +4.7 78.97 +6.1 88.72 –0.9 93.89 +0.4 93.78 -0.4
@@
TransMIL+PackMIL (Ours) 74.75 +6.2 78.87 +6.3 78.88 +7.0 92.31 +1.0 94.37 +0.3 94.12 +0.6
@@
RRTMIL+PackMIL (Ours) 74.63 +4.6 78.46 +6.3 78.43 +6.1 92.43 +0.9 94.54 +0.4 94.47 +0.4
@@ 4.2 Main Results
Comparison Methods. We compare several established and recent MIL aggregators Ilse et al. (2018), Lu et al. (2021), Shao et al. (2021), Li et al. (2021; 2024c), Tang et al. (2024), Zhang et al. (2024a; 2022) using three SOTA pathology encoders: UNI Chen et al. (2024), CHIEF Wang et al. (2024b), and GigaPath (GIGAP) Xu et al. (2024). To comprehensively evaluate the proposed framework, we select four widely-used MIL models as baselines. Additionally, we compare PackMIL against a standard random sampling (RS) strategy (i.e., sampling all inputs to a fixed length for batched training) to assess its effectiveness.
Table 2: Comparison of Survival Analysis (C-index Harrell Jr et al. (1996)) on BRCA and LUAD. OOM denotes Out-of-Memory in 24GB-RTX3090.
@@
ABMIL Ilse et al. (2018) 65.36±9.00 65.87±9.43 65.72±8.86 61.99±8.70 60.54±8.83 59.88±8.60
DSMIL Li et al. (2021) 65.81±9.17 65.87±9.98 64.94±9.22 62.38±8.50 61.44±8.76 61.35±8.20
CLAM Lu et al. (2021) 65.03±9.38 65.45±10.0 63.91±9.70 61.78±8.84 59.71±8.48 60.70±8.68
TransMIL Shao et al. (2021) 65.75±8.87 65.34±9.44 65.35±9.18 63.68±8.66 62.63±8.70 62.53±8.53
DTFD Zhang et al. (2022) 67.22±8.91 65.05±10.3 65.54±9.08 62.87±8.56 60.83±8.64 60.88±8.67
WIKG Li et al. (2024c) 65.55±9.14 65.77±9.21 65.79±9.62 OOM OOM OOM
RRTMIL Tang et al. (2024) 67.13±8.77 66.71±9.95 66.31±9.48 63.51±8.76 61.32±8.73 62.41±8.41
Mamba2D Zhang et al. (2024a) 66.01±7.06 65.73±8.15 65.96±7.69 63.23±6.70 60.69±6.97 62.19±6.73
ABMIL+RS 64.02 –1.3 65.71 –0.2 63.88 –1.8 61.98 -0.1 60.34 –0.2 61.01 +1.0
ABMIL+PackMIL (Ours) 68.30 +2.9 68.14 +2.3 67.04 +1.3 63.72 +1.8 62.60 +1.0 61.58 +1.7
DSMIL+RS 65.00 –0.8 64.72 –1.1 65.69 +0.8 63.28 +0.9 61.53 +0.1 61.00 –0.4
DSMIL+PackMIL (Ours) 69.76 +2.9 70.00 +4.1 68.03 +3.1 64.10 +1.7 62.18 +0.7 62.44 +1.1
@@
Focusing on Data Challenges in the FM Era. Although MIL architectures have advanced significantly, their performance improvements become marginal when using offline features extracted by foundation models (FMs). The quality of the FM primarily determines the final performance, and the latest or more complex MIL methods have reached a performance bottleneck. In this context, we observe that addressing the inherent data challenges in CPath is an effective way to achieve significant performance gains. Specifically, batched training based on random sampling (RS) achieves substantial improvements on grading tasks (Tab. 1), especially on large datasets such as PANDA ($\sim$10,000 slides). We attribute this primarily to the advantages of batched training on large-scale data and its effectiveness in reducing input redundancy. However, this RS strategy performs inconsistently on benchmarks like TCGA-BRCA-subtyping, which exhibits greater data heterogeneity (e.g., a sequence length variation of 60,000 compared to 1,000 in the PANDA dataset). It provides only marginal gains on some benchmarks while degrading performance on others in this task. As shown in Tab. 2, this issue becomes more pronounced in survival analysis, where sequence length variation are even larger. This sensitivity to heterogeneity is further reflected in its performance on complex methods like RRTMIL, where RS shows a significant performance gap across all benchmarks. These results indicate that RS compromises the WSI heterogeneity preserved during traditional training (batchsize = 1) and suffers from feature loss due to sampling.
PackMIL More Comprehensively Alleviates Data Challenges. Compared with RS, PackMIL more comprehensively alleviates the various data challenges in CPath, achieving more substantial and consistent performance gains. For example, its superior performance in survival analysis demonstrates its ability to handle data heterogeneity. Moreover, PackMIL yields significant further improvements even on grading tasks where RS already performed well. We attribute this gain to PackMIL’s ability to mitigate challenges of insufficient supervision. Notably, by comprehensively addressing these data challenges, PackMIL enables models to achieve performance comparable to higher-quality features with lower-cost offline features. Furthermore, it achieves consistent improvements when paired with better features, demonstrating its generalizability. In summary, results demonstrate the significant impact of inherent data challenges on CPath performance and validate the effectiveness of PackMIL.
@@ 4.3 Ablation Study
In this subsection, we systematically ablate the components of PackMIL. By default, we use ABMIL as the baseline model and UNI as the offline feature extractor. For the survival analysis task, we utilize the larger BRCA dataset. Appendix B gives extra discussions.
Batched Training and Packing Strategy. The results at the middle of Tab. 3 show that Random Sampling (RS) achieves significant improvements on the grading task, but its performance degrades on the other two tasks. This is attributed to its effective mitigation of input redundancy. However, this approach compromises data heterogeneity and leads to feature loss. Furthermore, batched training yields consistent performance improvements (row 7). Importantly, it leads to significant gains in training stability and efficiency. Specifically, while maintaining faster and more stable convergence (Tab. 3(a)), it also alleviates overfitting on the test set. With the pack-based batched training scheme, PackMIL retains data heterogeneity. This allows PackMIL to benefit from batch processing while preserving the performance advantages of traditional training (batchsize=1).
Table 3: Top: Ablation of PackMIL and computational cost on PANDA. TTime (RTX 3090 GPU-hours) denotes Train Time. Memory is the GPU memory which evaluated during training. FPS stands for frames per second. ADS module is disabled by default on PANDA; we detail its computational cost in Appendix B.3. Bottom: Loss curves of main and hyperslide loss under different settings.
@@
+ Random Sampling + HyperSlide $\checkmark$ 2.5h 2.5G 2056 79.93 94.37 67.15
@@
Hyperslide. We constructs a hyperslide using discarded features in the residual branch and task-relevant hyperslide labels, aiming to offer multi-slide supervision while mitigating feature loss from sampling. Tab. 3(b) confirms that the hyperslide learns effectively when guided by the proposed task-relevant labels. Furthermore, during joint training with the task loss, we find that optimizing the hyperslide also helps mitigate the rapid overfitting (i.e., the model rapid convergence of the training loss to near-zero on the training data within early epochs) of the model to the task loss on the FM features (Tab. 3(c)). This issue hinders the model from benefiting from the task loss, consequently reducing training quality. The results in row 8 and row 10 of Tab. 3 demonstrate the performance improvements achieved by multi-slide supervision and minimizing feature loss.
Attention-driven Downsampler. The Attention-driven Downsampler (ADS) is designed to mitigate potential input redundancy via feature fusion. Since the PANDA dataset contains fewer patches per WSI, we do not employ ADS on this task. However, for the sub-typing and survival tasks, which involve longer input sequences, the results indicate that ADS effectively reduces input redundancy. More important, through attention weighting and feature fusion, ADS preserves the discriminative information of the original features. Appendix B.3 gives further discussion.
@@
5 Conclusion

### v2

@@ 3.3 Training Recipe
Task-specific Hyperslide Labels and Loss. Heterogeneous pathological data and diverse downstream clinical tasks necessitate task-specific hyperslide labeling strategies that preserve task-relevant clinical characteristics and enable efficient learning from multiple WSIs. We divide downstream tasks into two classes: Morphological Categorical and Event-Driven. Based on pathological scenarios, we design three strategies to generate a hyperslide label $y^{\mathrm{hyper}}$ for higher-level supervision, illustrated in Fig.4. Appropriate loss functions are selected for different tasks, as detailed in the Supplementary.
1) Grading. Let each pack contain $S$ WSIs, each associated with a Gleason Score $g_{s}$ (determined by the area proportions of the primary and secondary Gleason patterns) and a patch count $n_{s}$. We compute a pack-level statistic by $\tilde{g}=\frac{\sum_{s=1}^{S}n_{s}\,g_{s}}{\sum_{s=1}^{S}n_{s}}$. The weighted metric is then mapped to a discrete ISUP grade $y^{\mathrm{hyper}}\in\{1,\dots,G\}$, serving as the single categorical hyperslide label. Weighting by $n_{s}$ ensures that the pack-level grading reflects the relative tissue coverage of each WSI, preventing smaller tissue regions from being overrepresented in the final assessment.
@@
where $\mathcal{E}$ is the ordered event set. This strategy ensures that each pack is uniquely associated with the most clinically significant event.
Figure 5: Practical Guidelines for Batched CPath Training. Impact of training hyperparameters, highlighting dataset-scale-dependent strategies. (a) On the large-scale PANDA dataset, accuracy is non-monotonic with bsbs, degrading at too large values. (b) On the smaller TCGA dataset, computational resource limitations necessitate an empirically tuned trade-off between bsbs and number of instances. (c) The $\sqrt{bs}$ learning rate scaling rule is effective for PANDA (bottom) but fails on TCGA (top), showing standard rules are not universal in CPath.
Practical Guidelines for Batched Training. Selecting an appropriate batchsize ($bs$) across various CPath tasks remains an underexplored yet critical problem for effective model training. To establish an batched training protocol, we conduct a series of preliminary experiments, as shown in Fig. 5. The results from these experiments inform the practical guidelines discussed below, which are subsequently adopted in our main experimental setup. We find that the specifics of these guidelines vary significantly across benchmarks and are strongly correlated with dataset scale.
On conventional datasets (e.g., TCGA), which typically contain a limited number of WSIs , resource constraints often force a trade-off between the $bs$ and the number of patches sampled per WSI (NP), as shown in Fig. 5(b). The optimal choice of $bs$ usually requires empirical tuning. Additionally, learning rate is not scaled by regular rules and also requires empirical tuning, as depicted in Fig. 5(c).
On large-scale datasets (e.g., PANDA [panda]), the $\sqrt{bs}$ learning rate scaling rule often proves effective. However, performance does not monotonically increase with $bs$; excessively large values can degrade performance, as illustrated in Fig. 5(a). Moreover, we found that 1D Batch Normalization [ioffe2015batch] can effectively facilitate convergence with sufficient data scale. In contrast, normalization does not provide significant improvements on conventional CPath datasets, which we attribute to their limited data volume.
@@ 4.1 Datasets and Evaluation Metrics
For cancer diagnosis, we evaluate performance on grading and sub-typing tasks using the PANDA [panda] and TCGA-BRCA datasets. For cancer prognosis, we evaluate survival analysis performance using TCGA-LUAD and TCGA-BRCA. We report macro accuracy (Acc.) for cancer grading and the area under the ROC curve (AUC) for sub-typing. For survival analysis, we report the concordance index (C-index) [c-index_harrell]. To ensure statistical robustness, we perform 1000 bootstrap iterations and repeat 5 experiments for grading and subtyping, and 9 times for survival. We report the mean and 95% confidence interval for all metrics. Supplementary also presents performance on the CAMELYON [c16] and fundus datasets.
Table 1: Comparison of Grading (Acc.) on PANDA and Sub-typing (AUC) on BRCA. Supplementary gives 95% CI of RS and PackMIL.
@@
ABMIL [ilse2018attention] 65.48±0.70 73.21±0.61 72.91±0.62 89.58±5.15 93.58±3.92 94.13±3.88
DSMIL [li2021dual] 71.95±0.72 71.41±0.68 70.84±0.73 90.68±4.91 93.89±3.79 93.70±4.14
CLAM [clam] 66.91±0.73 74.76±0.61 74.99±0.60 89.61±5.05 93.90±3.86 93.73±3.75
TransMIL [shao2021transmil] 68.53±0.79 72.59±0.75 71.91±0.74 91.91±4.34 94.09±3.79 93.57±3.84
DTFD [zhang2022dtfd] 63.57±0.71 72.19±0.66 71.93±0.65 91.07±4.59 93.86±3.93 93.88±3.81
WiKG [li2024dynamic] 72.37±0.72 76.68±0.68 74.91±0.68 91.93±4.65 93.81±4.25 94.06±3.92
RRTMIL [tang2024feature] 69.99±0.69 72.18±0.67 72.34±0.63 91.67±4.77 94.10±3.65 94.04±3.86
2DMamba [zhang20242dmamba] 71.59±0.73 74.97±0.68 75.36±0.66 90.96±3.85 93.59±3.14 93.33±3.25
ABMIL+RS 74.72 +9.2 77.91 +4.7 78.97 +6.1 88.72 –0.9 93.89 +0.3 93.78 -0.4
@@
TransMIL+PackMIL (Ours) 74.75 +6.2 78.87 +6.3 78.88 +7.0 92.31 +0.4 94.37 +0.3 94.12 +0.6
@@
RRTMIL+PackMIL (Ours) 74.63 +4.6 78.46 +6.3 78.43 +6.1 92.43 +0.8 94.54 +0.4 94.47 +0.4
@@ 4.2 Main Results
Comparison Methods. We compare several established and MIL aggregators [ilse2018attention, clam, shao2021transmil, li2021dual, li2024dynamic, tang2024feature, zhang20242dmamba, zhang2022dtfd] using three SOTA pathology encoders: UNI [uni], CHIEF [chief], and GigaPath (GIGAP) [gigap]. Supplementary gives the additional comparisons with CONCHv1.5 [lu2024avisionlanguage] and TITAN [ding2025multimodal]. To comprehensively evaluate the proposed framework, we select four widely-used MIL models as baselines. Additionally, we compare PackMIL against a standard random sampling (RS) strategy (i.e., sampling all inputs to a fixed length for batched training) to assess its effectiveness.
Table 2: Comparison of Survival Analysis (C-index [c-index_harrell]) on BRCA and LUAD. OOM denotes Out-of-Memory in 24GB-RTX3090.
@@
ABMIL [ilse2018attention] 65.36±9.00 65.87±9.43 65.72±8.86 61.99±8.70 60.54±8.83 59.88±8.60
DSMIL [li2021dual] 65.81±9.17 65.87±9.98 64.94±9.22 62.38±8.50 61.44±8.76 61.35±8.20
CLAM [clam] 65.03±9.38 65.45±10.0 63.91±9.70 61.78±8.84 59.71±8.48 60.70±8.68
TransMIL [shao2021transmil] 65.75±8.87 65.34±9.44 65.35±9.18 63.68±8.66 62.63±8.70 62.53±8.53
DTFD [zhang2022dtfd] 67.22±8.91 65.05±10.3 65.54±9.08 62.87±8.56 60.83±8.64 60.88±8.67
WIKG [li2024dynamic] 65.55±9.14 65.77±9.21 65.79±9.62 OOM OOM OOM
RRTMIL [tang2024feature] 67.13±8.77 66.71±9.95 66.31±9.48 63.51±8.76 61.32±8.73 62.41±8.41
2DMamba [zhang20242dmamba] 66.01±7.06 65.73±8.15 65.96±7.69 63.23±6.70 60.69±6.97 62.19±6.73
ABMIL+RS 64.02 –1.3 65.71 –0.2 63.88 –1.8 61.98 -0.0 60.34 –0.2 61.01 +1.1
ABMIL+PackMIL (Ours) 68.30 +2.9 68.14 +2.3 67.04 +1.3 63.72 +1.7 62.60 +2.1 61.58 +1.7
DSMIL+RS 65.00 –0.8 64.72 –1.2 65.69 +0.8 63.28 +0.9 61.53 +0.1 61.00 –0.4
DSMIL+PackMIL (Ours) 69.76 +4.0 70.00 +4.1 68.03 +3.1 64.10 +1.7 62.18 +0.7 62.44 +1.1
@@
Focusing on Data Challenges in the FM Era. Although MIL architectures have advanced significantly, their performance improvements become marginal when using offline features extracted by foundation models (FMs). For example, with UNI features, the gap between the latest and classic MIL methods in TCGA-BRCA-subtyping is only 0.52%. The quality of the FM primarily determines the final performance, and the latest or more complex MIL methods have reached a performance bottleneck [chen2024benchmarking, shaomultiple]. In this context, we observe that addressing the inherent data challenges in CPath is an effective way to achieve significant performance gains. Specifically, random sampling (RS) achieves substantial improvements on grading tasks (Tab. 1), especially on large datasets such as PANDA ($\sim$10,000 slides). We attribute this primarily to its effectiveness in reducing input redundancy and the advantages of batched training on large-scale data. However, this RS strategy performs inconsistently on benchmarks like TCGA-BRCA-subtyping, which exhibits greater data heterogeneity (e.g., a sequence length variation of 60,000 compared to 1,000 in the PANDA dataset). It provides only marginal gains on some benchmarks while degrading performance on others in this task. As shown in Tab. 2, this issue becomes more pronounced in survival analysis, where sequence length variation are even larger. This sensitivity to heterogeneity is further reflected in its performance on complex methods like RRTMIL, where RS shows a significant performance gap across all benchmarks. These results indicate that RS compromises the WSI heterogeneity preserved during traditional training (batchsize = 1) and suffers from feature loss due to sampling.
PackMIL More Comprehensively Alleviates Data Challenges. Compared with RS, PackMIL more comprehensively alleviates the various data challenges in CPath, achieving more substantial and consistent performance gains (+1.28% on TCGA-BRCA-subtyping). For example, its superior performance in survival analysis demonstrates its ability to handle data heterogeneity. Moreover, PackMIL yields significant further improvements even on grading tasks where RS already performed well. We attribute this gain to PackMIL’s ability to mitigate challenges of insufficient supervision. Notably, by comprehensively addressing these data challenges, PackMIL enables models to achieve performance comparable to higher-quality features with lower-cost offline features. Furthermore, it achieves consistent improvements with better features, demonstrating its generalizability. In summary, results demonstrate the significant impact of data challenges on CPath performance and validate the effectiveness of PackMIL.
@@ 4.3 Ablation Study
Unless otherwise specified, our ablations use ABMIL as the baseline model and UNI as the offline feature extractor. For the survival analysis task, we utilize the larger BRCA dataset. Supplementary gives further discussion for each subsection, qualitative analysis, inference efficiency and key hyperparameters (i.e., $k$, $\lambda$, and branch split ratio).
Table 3: Top: Ablation of PackMIL and computational cost on PANDA. TTime (RTX 3090 GPU-hours) denotes Train Time. Memory is the GPU memory which evaluated during training. FPS stands for frames per second. ADS module is disabled by default on PANDA; we detail its computational cost in Supplementary. Bottom: Loss curves of main and hyperslide loss under different settings.
@@
+ Random Sampling + HyperSlide $\checkmark$ 2.5h 2.5G 2056 79.93 94.02 67.15
@@
Batched Training and Packing Strategy. The results at the middle of Tab. 3 show that Random Sampling (RS) achieves significant improvements on the grading task, but its performance degrades on the other two tasks. This is attributed to its effective mitigation of input redundancy. However, this approach compromises data heterogeneity and leads to feature loss. Furthermore, batched training yields consistent performance improvements (row 6). Importantly, it leads to significant gains in training stability and efficiency. Specifically, while maintaining faster and more stable convergence (Tab. 3(a)), it also alleviates overfitting on the test set. With the pack-based batched training, PackMIL retains data heterogeneity. This allows PackMIL to benefit from batch processing while preserving the performance advantages of traditional training (batchsize=1).
Hyperslide. We constructs a hyperslide using discarded features in the residual branch and task-relevant hyperslide labels, aiming to offer multi-slide supervision while mitigating feature loss from sampling. Tab. 3(b) confirms that the hyperslide learns effectively when guided by the proposed task-relevant labels. Furthermore, during joint training with the task loss, we find that optimizing the hyperslide also helps mitigate the rapid overfitting (i.e., the model rapid convergence of the training loss to near-zero on the training data within early epochs) of the model to the task loss on the FM features (Tab. 3(c)). This issue hinders the model from benefiting from the task loss, consequently reducing training quality. The results in row 7 and row 9 of Tab. 3 demonstrate the performance improvements achieved by multi-slide supervision and minimizing feature loss.
Attention-driven Downsampler. The Attention-driven Downsampler (ADS) is designed to mitigate potential input redundancy via feature fusion. Since the PANDA dataset contains fewer patches per WSI, we do not employ ADS on this task. However, for the sub-typing and survival tasks, which involve longer input sequences, the results indicate that ADS effectively reduces input redundancy. More important, through attention weighting and feature fusion, ADS preserves the discriminative information of the original features. Supplementary gives further discussion.
@@
5 Discussion
Why Packing. Packing enables memory-efficient batched training while preserving the heterogeneity of WSI data. When dealing with significant sequence length variations (from 200 to 200k), padding is memory-inefficient, while sampling compromises data heterogeneity.
Why Hyperslide. While constructing supervision labels within a single bag (intra-slide) is a mature approach in this field, Hyperslide packs instances from different bags to enable inter-slide supervision and knowledge supplementation. We believe this is a pioneering attempt that aligns with the clinical experience of pathologists.
Packing Any MIL. PackMIL can be adapted to arbitrary MIL models. We have already implemented support for the most common and basic MIL architectures. Even if adapting certain complex architectures proves challenging, they can still be trained in a non-parallel fashion.